CONFORMATIONAL ISOMERIZATION KINETICS OF VINYL ISOCYANATE MEASURED BY DYNAMIC ROTATIONAL SPECTROSCOPY

\underline{JUSTIN L. NEILL}, MATT T. MUCKLE, and BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904.

Vinyl isocyanate (CH_{2}=CH–N=C=O) can exist in two stable conformations, having a cis or trans arrangement about the C-N single bond. Electronic structure calculations indicate that the trans conformer is about 250 cm^{−1} lower in energy. There is a low barrier to isomerization of the trans conformer calculated to be about 580 cm^{−1}. These energetics are supported by the pure rotational spectrum of vinyl isocyanate in a molecular beam, where transitions of the trans conformer are about 300 times stronger than those of the cis conformer due to conformational relaxation in the freejet expansion. The dominance of the trans conformer guarantees that vibrational excited states prepared by pulsed infrared laser excitation are effectively conformer-selective. Dynamic rotational spectra of laser-prepared excited states in the 3000 cm^{−1} region of the spectrum show evidence of conformational isomerization through coalescence of the overall line shape. Strong mode-specific reaction yields are observed from different vibrational bands where, in some frequency regions, there is no evidence of conformational isomerization despite the molecule having a total vibrational energy greatly exceeding the barrier to reaction. The nuclear quadrupole hyperfine structure of the dynamic rotational spectra shows that isomerization, when it occurs, conserves the K_{a} angular momentum projection quantum number.